# SenseSeek Dataset: Multimodal Sensing to Study Information Seeking Behaviors

### v1

@@ Abstract.
Information processing tasks involve complex cognitive mechanisms that are shaped by various factors, including individual goals, prior experience, and system environments. Understanding such behaviors requires a sophisticated and personalized data capture of how one interacts with modern information systems (e.g., web search engines). Passive sensors, such as wearables, capturing physiological and behavioral data, have the potential to provide solutions in this context. This paper presents a novel dataset, SenseSeek, designed to evaluate the effectiveness of consumer-grade sensors in a complex information processing scenario: searching via systems (e.g., search engines), one of the common strategies users employ for information seeking. The SenseSeek dataset comprises data collected from 20 participants, 235 trials of the stimulated search process, 940 phases of stages in the search process, including the realization of Information Need (IN), Query Formulation (QF), Query Submission by Typing (QS-T) or Speaking (QS-S), and Relevance Judgment by Reading (RJ-R) or Listening (RJ-L). The data includes Electrodermal Activities (EDA), Electroencephalogram (EEG), PUPIL, GAZE, and MOTION data, which were captured using consumer-grade sensors. It also contains 258 features extracted from the sensor data, the gaze-annotated screen recordings, and task responses. We validate the usefulness of the dataset by providing baseline analysis on the impacts of different cognitive intents and interaction modalities on the sensor data, and effectiveness of the data in discriminating the search stages. To our knowledge, SenseSeek is the first dataset that characterizes the multiple stages involved in information seeking with physiological signals collected from multiple sensors. We hope this dataset can serve as a reference for future research on information-seeking behaviors.
@@ 1. Introduction
Information is ubiquitous (Fidel, 2012), and modern information access systems such as web search engines or intelligent assistants have evolved to become increasingly diverse, offering greater mobility and multiple ways of interaction. In this context, wearable sensors that capture physiological and behavioral data have demonstrated remarkable capabilities in enhancing human-computer interaction and improving daily life applications. There is growing research interest in leveraging these wearable sensors to explore novel questions in information interaction, including the detection of user interest (Arapakis et al., 2019; Edwards and Kelly, 2017) and engagement during video browsing (He et al., 2023; Banik et al., 2024), recognition of cognitive preference (Boonprakong et al., 2023; Moravec et al., 2018) and emotional responses (Gebhardt et al., 2024) on social media content, and understanding various information needs (Michalkova et al., 2022).
@@ 4.3. Data Pre-Processing
All data was processed with the Python programming language. EEG was processed using the MNE library  [^5] https://mne.tools/stable/i. EDA and wrist motion data obtained from the wristband were processed using the NeuroKit2 (Makowski et al., 2021) library. PUPIL data was cleaned with the PyPlr (Martin et al., 2021) library. The data collected during the break section was excluded from pre-processing.
@@
Following similar procedures to Martínez-Santiago et al. (2023), Ye et al. (2024), Gwizdka et al. (2017), EEG data was first re-referenced with the common average, denoised with a Butterworth filter (1 – 40Hz), and removed the signal mean for the middle 90% percentile. Next, to remove the artifacts (e.g., blinking), the data was further cleaned and interpolated with the Autoreject (Jas et al., 2017) package and Independent Component Analysis (ICA) combined with ICLabel (Li et al., 2022). The power spectral density (PSD) of each EEG channel was then calculated using Welch’s method and hamming window and normalized (Kosonogov et al., 2023; So et al., 2017; Lee et al., 2020), for 4 frequency bands, $Theta$ (4–8Hz), $Alpha$ (8–13 Hz), $Beta$ (13–30Hz) and $LowGamma$ (25–40Hz) (Gwizdka et al., 2017).
@@ 5. Validation
Before conducting any analysis, we first checked the self-reported task perceptions to ensure the experimental conditions meet our anticipation (Section 5.1). To answer RQ1 and RQ2, we conducted the exploratory analysis with statistical tests in a within-subject setting. We first conducted non-parametric Friedman tests to compare across EYEOPEN baseline, and all search stages (IN, QF, QS-T, QS-S, RJ-R, RJ-L) for overall impacts (Section 5.2).
@@ 5.5.1. Experimental Setting
We conducted Leave-One-participant-Out (LOO) cross-validation and reported accuracy and F1-macro metrics for the evaluation. LOO cross-validation assesses the performance of a model by iteratively training on data from all participants except one, which is reserved for testing. In this project, which involves data from 20 participants, each model was evaluated 20 times, with the data from 19 participants used for training and the data from 1 participant used for testing each time [^6] Note that as 2 participants are missing the eye-tracking data, the models involve PUPIL or GAZE were evaluated 18 times.. Accuracy reflects the overall correctness of a model’s predictions, while the F1-macro score addresses class imbalance by averaging performance equally across all classes, offering a balanced evaluation of model effectiveness. Higher scores for both metrics indicate better model performance. We employed the random guess baseline, which is around 17%.
@@ 5.5.2. Single-Sensor Performance
Figure 10 reveals distinct classification patterns across models. The EDA model showed poor discriminative capability with predictions skewed toward IN and QS-T. While EEG provided more balanced predictions, it still favored QS-T and RJ-L. Only the PUPIL model achieved well-balanced and accurate predictions across all classes, while GAZE and MOTION models exhibited biases similar to EDA. Notably, when comparing the eye-driven activities IN and RJ-R (which involve different cognitive intentions), only MOTION and PUPIL models successfully differentiated between them. This suggests that similar gaze movement patterns in reading activities may confound the GAZE model, while the wristband may lack sufficient sensitivity to capture subtle EDA variations during these activities.
@@ 6. Discussion & Limitation
Considering the possible cognitive activities participants might engage in during these stages, at EYEOPEN, participants only need to focus on the screen and relax. At IN, participants may engage in cognitive processes such as comprehending information, retrieving memories, assessing personal interest, and updating knowledge gaps, which drive subsequent actions (Michalkova et al., 2022; Moshfeghi and Pollick, 2018; Michalkova et al., 2022). At RJ-R, beyond comprehension, participants also evaluated information based on factors such as usefulness and topical relevance, ultimately deciding whether to accept it (Allegretti et al., 2015; Ji et al., 2024; Moshfeghi and Pollick, 2018). Taken together, and in alignment with existing findings, our results indicated the lack of significant difference between IN and RJ-R implies these two tasks engage similar neural mechanisms, while EYEOPEN appears to invoke distinct attentional shifts.
@@
In this dataset, the lengths of the queries submitted by participants between QS-T and QS-S were similar. The text queries had an average of 8 words, ranging from 2 to 19 words, and the voice queries had an average of 7.8 words, ranging from 2 to 20 words. However, the interaction durations differed. Typing required more time (average: 17.7 seconds) compared to speaking (average: 6.8 seconds). These differences in interaction duration and underlying cognitive processing might manifest in changing physiological data (Iadarola et al., 2021). Interestingly, no significant differences were observed in any brain regions or peripheral physiological data between QS-T and QS-S. This might suggest that the modality to express the query – whether typed or spoken – may not significantly affect cognitive load or neural engagement captured by the physiological data.
@@ 6.4. Limitations & Ecological Validity
Moreover, search stage durations varied considerably, from 6.8 to 66.0 seconds on average, highlighting the dynamic nature of information searching and the importance of context-specific factors. This variability likely affected physiological responses, as longer stages may elicit more complex responses than shorter ones. On the other side, unlike most existing physiological datasets that typically involve longer engagement periods, this duration variability could benefit studies featuring rapid cognitive transitions of short-term interactions that characterize real-world search interactions.
Lastly, our study employed only consumer-grade sensors that offer lower accuracy and sampling rates than medical-grade equipment; hence, there is a possibility of artifacts or missed responses that may have impacted the results. As a trade-off, these sensors are closer to the actual technologies users encounter in everyday settings. Future research could bridge this gap by implementing a dual-monitoring approach, using both consumer- and medical-grade or more advanced equipment simultaneously to validate findings while maintaining ecological validity. Additionally, larger sample sizes would strengthen statistical analysis and help distinguish true physiological patterns. The advantages of the consumer-grade sensors, for example, accessibility, comfort, and ease of use, allow for a scalable version of this dataset.
@@ 7. Conclusion and Future Work
We presented SenseSeek, a comprehensive dataset examining the ramifications of passive sensing data and information activity, particularly focusing on the information search process. The dataset comprises physiological and behavioral data from 20 participants across 235 trials, 940 search stages in a controlled lab environment, where each participant completed simulated search tasks. Both the raw and cleaned data are made publicly available, along with 258 extracted features. To provide contextual richness, it also contains the task materials, self-rated perceptions, and gaze-annotated screen recordings. We validate the usefulness of the dataset by investigating how cognitive intents and interaction modalities affect information searching stages and evaluated the effectiveness of sensor data in distinguishing these stages. Statistically significant differences were observed in EEG and PUPIL data, particularly influenced by cognitive intents and interaction modality (reading vs. listening). Search stages classification was performed using features from individual or multiple sensor groups. Our analysis showed that PUPIL features performed best, and EDA features performed worst. The varying effectiveness of sensor groups underscores the need for sensor selection based on specific cognitive processes and modality-specific processing configurations to optimize classification performance in multi-modality models.
@@
By making this dataset publicly available and enhancing our understanding of these variations, we hope to contribute to the development of passive sensors and multimedia information access systems. For instance, designing adaptive search interfaces that respond to users’ cognitive needs in real time. This dataset lays a foundation for future research exploring more complex search activities and provides a rich source of data for developing algorithms that monitor user experiences such as attention, workload, and stress during information interaction. Researchers can leverage this baseline data to train computational models before applying them in naturalistic environments, ensuring more reliable transitions from controlled to real-world applications.
@@ 8. Dataset and Code Availability
The use of the SenseSeek dataset is limited to academic research purposes. The SenseSeek dataset is publicly available and can be accessed on the OSF platform at https://osf.io/waunb/?view_only=94756f9d2c7a49e094ae42d494c9516a. Additional descriptions of the dataset are included in the appendices. The code used for data processing, analysis, and model training is also available at https://github.com/kkkkk2017/SenseSeek-Dataset-code. The repository contains the instructions for the dataset, several Jupyter Notebook files with data analysis and visualizations, and Python source files for model training. All required packages are listed in the requirements.txt file.
@@ Acknowledgements.
This research is partially supported by the Sponsor Australian Research Council https://www.arc.gov.au/ (Grant #DE200100064, Grant #CE200100005).

### v2

@@ Abstract.
Information processing tasks involve complex cognitive mechanisms that are shaped by various factors, including individual goals, prior experience, and system environments. Understanding such behaviors requires a sophisticated and personalized data capture of how one interacts with modern information systems (e.g., web search engines). Passive sensors, such as wearables, capturing physiological and behavioral data, have the potential to provide solutions in this context. This paper presents a novel dataset, SenseSeek, designed to evaluate the effectiveness of consumer-grade sensors in a complex information processing scenario: searching via systems (e.g., search engines), one of the common strategies users employ for information seeking. The SenseSeek dataset comprises data collected from 20 participants, 235 trials of the stimulated search process, 940 phases of stages in the search process, including the realization of Information Need (IN), Query Formulation (QF), Query Submission by Typing (QS-T) or Speaking (QS-S), and Relevance Judgment by Reading (RJ-R) or Listening (RJ-L). The data includes Electrodermal Activities (EDA), Electroencephalogram (EEG), PUPIL, GAZE, and MOTION data, which were captured using consumer-grade sensors. It also contains 258 features extracted from the sensor data, the gaze-annotated screen recordings, and task responses. We validate the usefulness of the dataset by providing baseline analyses on the impacts of different cognitive intents and interaction modalities on the sensor data, and effectiveness of the data in discriminating the search stages. To our knowledge, SenseSeek is the first dataset that characterizes multiple stages involved in information seeking with physiological signals collected from multiple sensors. We hope this dataset can serve as a reference for future research on information-seeking behaviors.
@@ 1. Introduction
Information is ubiquitous (Fidel, 2012), and modern information access systems such as web search engines or intelligent assistants have evolved to become increasingly diverse, offering greater mobility and multiple ways of interaction. In this context, wearable sensors that capture physiological and behavioral data have demonstrated remarkable capabilities in enhancing human-computer interaction and improving daily life applications. There is growing research interest in leveraging these wearable sensors to explore novel questions in information interaction, including the detection of user interest (Arapakis et al., 2019; Edwards and Kelly, 2017) and engagement during video browsing (He et al., 2023; Banik et al., 2024), recognition of cognitive preference (Boonprakong et al., 2023; Moravec et al., 2019) and emotional responses (Gebhardt et al., 2024) on social media content, and understanding various information needs (Michalkova et al., 2022).
@@ 3.1. Participants
[…]
A flowchart illustrating the sequence of tasks participants completed during the study. It includes baseline recordings, topic presentation, query formulation, submission (typing or speaking), relevance judgment (reading or listening), and ratings. The diagram highlights randomized topic order and interaction modalities.
@@ 4.2. Sensor Synchronization
[…]
A timeline diagram showing the synchronization of sensor data collection during different phases of the experiment, including calibration, baseline, task execution, and breaks.
@@ 4.3. Data Pre-Processing
All data was processed with the Python programming language. EEG data were processed using the MNE library  [^5] https://mne.tools/stable/i. EDA and wrist motion data obtained from the wristband were processed using the NeuroKit2 (Makowski et al., 2021) library. PUPIL data was cleaned with the PyPlr (Martin et al., 2021) library. Data collected during the break section were excluded from pre-processing.
@@
Following similar procedures to Martínez-Santiago et al. (2023), Ye et al. (2024), Gwizdka et al. (2017), EEG data were first re-referenced with the common average, denoised with a Butterworth filter (1 – 40Hz), and removed the signal mean for the middle 90% percentile. Next, to remove the artifacts (e.g., blinking), the data was further cleaned and interpolated with the Autoreject (Jas et al., 2017) package and Independent Component Analysis (ICA) combined with ICLabel (Li et al., 2022). The power spectral density (PSD) of each EEG channel was then calculated using Welch’s method and hamming window and normalized (Kosonogov et al., 2023; So et al., 2017; Lee et al., 2020), for 4 frequency bands, $Theta$ (4–8Hz), $Alpha$ (8–13 Hz), $Beta$ (13–30Hz) and $LowGamma$ (25–40Hz) (Gwizdka et al., 2017).
@@ 5. Validation
Before conducting any analysis, we first reviewed the self-reported task perceptions to ensure the experimental conditions met our anticipation (Section 5.1). To answer RQ1 and RQ2, we conducted the exploratory analysis with statistical tests in a within-subject setting. We first conducted non-parametric Friedman tests to compare across EYEOPEN baseline, and all search stages (IN, QF, QS-T, QS-S, RJ-R, RJ-L) for overall impacts (Section 5.2).
@@ 5.1. Task Completeness & Manipulation Check
[…]
A bar chart displaying participants’ ratings of topic difficulty, familiarity, interest, and perceived relevance and difficulty of the search results. It confirms that the stimuli met the expected criteria.
@@ 5.2. Understanding Physiological Responses at Different Stages of Information Search
Topography heatmap showing EEG channel-wise statistical differences across search stages for four frequency bands (Theta, Alpha, Beta, Low Gamma). Highlighted regions indicate significant cognitive activity variations.
[…]
@@ B QS-S QS-T RJ-R RJ-L QS-T RJ-R RJ-L RJ-R RJ-L RJ-L
[…]
A violin plot comparing mean pupil dilation across different search stages and baseline. It visualizes cognitive load and attention levels during each stage.
@@ 5.3. The influence of Cognitive Intents: Comparing Eye-Driven Activities
(a) EEG topography showing differences in Alpha band activity across EYEOPEN, IN, and RJ-R stages. (b) Line graph showing changes in pupil dilation over time during these stages.
[…]
@@ 5.4. The Influence of Interaction Modalities
[…]
Topographic maps comparing EEG responses between typing vs. speaking (QS-T vs. QS-S) and reading vs. listening (RJ-R vs. RJ-L). Significant differences are marked.
@@ 5.5.1. Experimental Setting
We conducted Leave-One-participant-Out (LOO) cross-validation and reported accuracy and F1-macro metrics for the evaluation. LOO cross-validation assesses the performance of a model by iteratively training on data from all participants except one, which is reserved for testing. In this project, which involves data from 20 participants, each model was evaluated 20 times, with the data from 19 participants used for training and the data from 1 participant used for testing each time [^6] Note that as 2 participants are missing the eye-tracking data, the models involving PUPIL or GAZE were evaluated 18 times.. Accuracy reflects the overall correctness of a model’s predictions, while the F1-macro score addresses class imbalance by averaging performance equally across all classes, offering a balanced evaluation of model effectiveness. Higher scores for both metrics indicate better model performance. We employed the random guess baseline, which is around 17%.
@@ 5.5.2. Single-Sensor Performance
Figure 10 reveals distinct classification patterns across models. The EDA model showed poor discriminative capability with predictions skewed toward IN and QS-T. While EEG provided more balanced predictions, it still favored QS-T and RJ-L. Only the PUPIL model produced well-balanced and accurate predictions across all classes, while GAZE and MOTION models exhibited biases similar to EDA. Notably, when comparing the eye-driven activities IN and RJ-R (which involve different cognitive intentions), only MOTION and PUPIL models successfully differentiated between them. This suggests that similar gaze movement patterns in reading activities may confound the GAZE model, while the wristband may lack sufficient sensitivity to capture subtle EDA variations during these activities.
@@
[…]
A bar chart comparing classification performance (F1-macro) of models using different sensor types and time windows. PUPIL and GAZE show the highest scores.
@@
[…]
Confusion matrices for each sensor type showing how well each model classified the six search stages. PUPIL and MOTION models show more balanced predictions.
@@ 6. Discussion & Limitation
Considering the possible cognitive activities participants might engage in during these stages, at EYEOPEN, participants only need to focus on the screen and relax. At IN, participants may engage in cognitive processes such as comprehending information, retrieving memories, assessing personal interest, and updating knowledge gaps, which drive subsequent actions (Michalkova et al., 2022; Moshfeghi and Pollick, 2018). At RJ-R, beyond comprehension, participants also evaluated information based on factors such as usefulness and topical relevance, ultimately deciding whether to accept it (Allegretti et al., 2015; Ji et al., 2024; Moshfeghi and Pollick, 2018). Taken together, and in alignment with existing findings, our results indicated the lack of significant difference between IN and RJ-R implies these two tasks engage similar neural mechanisms, while EYEOPEN appears to invoke distinct attentional shifts.
@@
In this dataset, the lengths of the queries submitted by participants between QS-T and QS-S were similar. The text queries had an average of 8 words, ranging from 2 to 19 words, and the voice queries had an average of 7.8 words, ranging from 2 to 20 words. However, the interaction durations differed. Typing required more time (average: 17.7 seconds) compared to speaking (average: 6.8 seconds). These differences in interaction duration and underlying cognitive processing might manifest in changing physiological data (Iadarola et al., 2021). Interestingly, no significant differences were observed in any brain regions or peripheral physiological data between QS-T and QS-S. This suggests that the modality used to express the query – typed or spoken – may not significantly affect cognitive load or neural engagement captured by the physiological data.
@@ 6.4. Limitations & Ecological Validity
Moreover, search stage durations varied considerably, from 6.8 to 66.0 seconds on average, highlighting the dynamic nature of information searching and the importance of context-specific factors. This variability likely affected physiological responses, as longer stages may elicit more complex responses than shorter ones. In contrast to most existing physiological datasets, which typically involve longer engagement periods, this duration variability could benefit studies featuring rapid cognitive transitions of short-term interactions that characterize real-world search interactions.
Lastly, our study employed only consumer-grade sensors that offer lower accuracy and sampling rates than medical-grade equipment; hence, there is a possibility of artifacts or missed responses that may have impacted the results. As a trade-off, these sensors are closer to the actual technologies users encounter in everyday settings. Future research could bridge this gap by implementing a dual-monitoring approach, using both consumer- and medical-grade or more advanced equipment simultaneously to validate findings while maintaining ecological validity. Additionally, larger sample sizes would strengthen statistical analysis and help distinguish true physiological patterns. Advantages of consumer-grade sensors, such as accessibility, comfort, and ease of use, allow for a scalable version of this dataset.
@@ 7. Conclusion and Future Work
We presented SenseSeek, a comprehensive dataset examining the ramifications of passive sensing data and information activity, particularly focusing on the information search process. The dataset comprises physiological and behavioral data from 20 participants across 235 trials, 940 search stages in a controlled lab environment, where each participant completed simulated search tasks. Both the raw and cleaned data are made publicly available, along with 258 extracted features. To provide contextual richness, it also contains the task materials, self-rated perceptions, and gaze-annotated screen recordings. We validate the usefulness of the dataset by investigating how cognitive intents and interaction modalities affect information searching stages and evaluating the effectiveness of sensor data in distinguishing these stages. Statistically significant differences were observed in EEG and PUPIL data, particularly influenced by cognitive intents and interaction modality (reading vs. listening). Search stages classification was performed using features from individual or multiple sensor groups. Our analysis showed that PUPIL features performed best, and EDA features performed worst. The varying effectiveness of sensor groups underscores the need for sensor selection based on specific cognitive processes and modality-specific processing configurations to optimize classification performance in multi-modality models.
@@
By making this dataset publicly available and deepening our understanding of these variations, we hope to contribute to the development of passive sensors and multimedia information access systems. For instance, designing adaptive search interfaces that respond to users’ cognitive needs in real-time. This dataset lays a foundation for future research exploring more complex search activities and provides a rich source of data for developing algorithms that monitor user experiences such as attention, workload, and stress during information interaction. Researchers can leverage this baseline data to train computational models before applying them in naturalistic environments, ensuring more reliable transitions from controlled to real-world applications.
@@ 8. Dataset and Code Availability
The use of the SenseSeek dataset is limited to academic research purposes. The SenseSeek dataset is publicly available and can be accessed on the OSF platform at https://osf.io/waunb/?view_only=94756f9d2c7a49e094ae42d494c9516a. Additional descriptions of the dataset are included in the appendices. The code used for data processing, analysis, and model training is also available at https://github.com/ADMSCentre/SenseSeek-Dataset-code. The repository contains the instructions for the dataset, several Jupyter Notebook files with data analysis and visualizations, and Python source files for model training. All required packages are listed in the requirements.txt file.
@@ Acknowledgements.
This research is partially supported by the Sponsor Australian Research Council https://www.arc.gov.au/ and the Sponsor Australian Research Council Centre of Excellence for Automated Decision-Making and Society https://www.admscentre.org.au/ (Grant #DE200100064, Grant #CE200100005).
@@ Appendix A Additional Description of the SenseSeek dataset
[…]
Screenshots from the experiment showing gaze tracking overlays on the interface. These visuals demonstrate how participants visually engaged with the content.